# Synthesis and stability of biomolecules under Earth’s upper mantle conditions

### v1

@@ II Results and Discussion
We first consider Ribose (C5H10O5), composed solely of C, H, O elements. The formose reaction is a well-known prebiotic pathway for synthesizing ribose from formaldehyde (CH2O) breslow1959mechanism . Our simulations have generated a few relevant hydrocarbon products (shown in Figures S8 and S9), including formaldehyde molecule (number 1 in Figure S8), so we chose it as the primary reactant of synthesizing ribose. Additionally, to facilitate the synthesis process, the reactants also include a “C2” molecule, (Z)-ethene-1,2-diol (C2H4O2, number 7 in Figure S8). Initially, We constructed a simulation box consisting of 10 (Z)-ethene-1,2-diol molecules and 5 formaldehyde molecules (Figure S10). After about 76 ps, a “C5” molecule was formed at $\sim$ 10 GPa and 1400 K. However, it could not transform into the structure of ribose after an extended time. We further inserted 5 water molecules to the system, and finally got ribose molecules. The reaction process is shown in Fig. 5 (a), where formaldehyde and “C2” molecules reacted to form a “C3” molecule, which subsequently combined with another “C2” molecule to form a “C5” molecule. This “C5” molecule then evolved into the open-chain form of ribose at about 19 ps.
@@
To analyze the formation process, we calculated the free energy in Fig. 5 (b), where two C-C distances were selected as collective variables labelled as CV1 and CV2. For the first collective variable, two carbon atoms come from formaldehyde and “C2” molecule, respectively, which are labeled as C1 with yellow color. While for the second collective variable, two carbon atoms are from two different “C2” molecules, marked as C2 with orange color. Our aim is to determine whether ribose formation follows the C1-C1-C2-C2 connectivity pattern. The free energy landscape shows a clear energy minimum that corresponds to the formation of ribose. However, the C-C distance at this point is larger than the typical C-C bond distance, indicating that the formation of ribose deviates from our initial expectations. To gain further insights, we examined the biased molecular dynamics trajectories and found that the carbon backbone of generated ribose molecule follows the C1-C2-C1-C2 arrangement. By measuring the distances of C1-C1 and C2-C2 in this molecule, we confirm that they correspond to the distances observed at the minimum energy point. Furthermore, an alternative initial configuration for the free energy calculation is also presented in Figure S11. Despite the unexpected bondings of the carbon atoms, the free energy landscape remains a reliable indicator of the ribose formation process, as the region of minimum energy still corresponds to a formation of ribose molecule.
@@
Previous studies showed that the cyclic forms of ribose are abundant in aqueous solutions at room temperature drew199813c . However, the cyclic forms of ribose were not detected after 90 ps in our simulations (Fig. 5 (a)). We also carried out further simulations to try to generate rings, as illustrated in Figure S12, but there was still no ring formed. However, that does not mean that the cyclic forms of ribose could not be formed under extreme conditions. The simulation time of 90 ps, though relatively long for AIMD simulations, is significantly shorter than the timescales of real chemical evolutions in deep Earth. The spacial limitation under high pressure may restrict the configuration changes of ribose, reducing the probability of the C4’ or C5’ hydroxyl group attacking the C1’ aldehyde group, which is disadvantageous for producing the furanose form (5-membered ring) or pyranose form (6-membered ring). These two factors suggest that some cyclic formation reactions may be restricted by the shorter timescale and small size of the system in AIMD simulations. To better study the cyclic formation mechanism, we chose the curved open-chain ribose molecules obtained from the previous simulations, as it appears that the C4’ or C5’ hydroxyl group has a greater chance of reacting with C1’ aldehyde group under such narrow space. We built a new simulation box containing three of these ribose molecules and eighteen water molecules. After about 36ps, a furanose form of ribose was formed. Fig. 5 (c) shows the ring formation mechanism, the C4’ hydroxyl group first lost a hydrogen atom, leaving a carbonyl group. The oxygen atom then attacked C1’ to form the ring, and finally, the carbonyl group on C1’ received one dissolved proton in the solution to generate the 5-membered ribose ring. No 6-membered ring was found from our simulations of extreme P-T conditions.
@@
We further applied free energy calculations to study the formation ability of 5- and 6-membered ribose rings at $\sim$ 10 GPa and 1400 K. We chose two C-O distances as the CVs for the two rings, as shown in Fig. 5 (d). The energy minima are at the C-O distance of 1.37 Å for the two ribose forms, indicating the formation of C-O bonds and the rings. Besides, the formation of the C-O bonds lowers the free energy by 11.57 kcal/mol for the 5-membered ribose ring and 5.85 kcal/mol for the 6-membered ribose ring , indicating that the ring formation is thermodynamically more stable and the furanose form of ribose is more favored than the pyranose form at $\sim$ 10 GPa and 1400 K. The free energy result explained why the the pyranose form of ribose is absent in our AIMD simulations. Compared with ambient conditions, high pressure would induce the limited space, which may not support the existence or formation of bigger cyclic form of molecules, so energy barrier for forming 5-membered ring is lower than that of 6-membered ring.
@@
Note that at ambient conditions, the naturally-occurring ribose exists as a mixture of cyclic forms in equilibrium with its open-chain form in aqueous solutions, and the pyranose form of ribose is the predominant conformation, accounting for approximately 70% dewick2013essentials ; dass2021equilibrium . However, despite the thermodynamic stability of the pyranose form of ribose over the furanose form at ambient conditions, the exclusive constituent of the carbohydrate backbone of RNA remains the furanose form of ribose. The reason behind it is still unknown. There are several hypotheses proposed to explain this preference, including silicate(borate)/ribose complexes kolb2004complexes ; lambert2004silicate and temperature gradients dass2021equilibrium . Our finding suggests that extreme P-T conditions in Earth’s upper mantle may also generate the necessary component for assembling RNA molecules. Consequently, it highlights the potential role of upper mantle environment in facilitating the emergence and development of RNA-based life.
@@
We then shifted our attention towards the chemical evolution of nucleobases, which are another essential building blocks of genetic information. There are five naturally-occurring nucleobases: adenine (A), cytosine (C), guanine (G), thymine (T), and uracil (U). Of these, uracil is unique to RNA, so our studies are primarily focused on investigating the formation of uracil molecules under extreme conditions. We first studied the formation of the open-chain “C4N2” molecule with the same CN-backbone present in uracil, as illustrated in Fig. 6 (a). Previous experimental studies on the origins of life have suggested that formamide is a crucial molecule for the prebiotic formation of nucleobases, in the presence of mineral or metal oxide catalysts saladino2012formamide , heatingenchev2021chemical , and high-energy free radicals ferus2015high ; ferus2014high , so here we also first chose formamide as a reactant. However, after about 144 ps, we did not observe the formation of the “C4N2” molecule at 10 GPa and 1400 K, indicating that it is difficult for small molecules to directly synthesize uracil in the AIMD simulations at extreme P-T conditions (Figure S13). Later, we chose the chemical reactants that have closer structural similarities to the “C4N2” molecule. The reactant molecules, including “C4N”, “C3N2”, “C2N2”, and “C3”, were generated from our previous AIMD simulations. One previous study suggested that uracil could be synthesized from the reactions between urea and “C3” molecules choe2022mechanism , so we specifically selected the C3H4O4 molecule (number 10 in Figure S8) and the urea molecule (CH4N2O, number 11 in Figure 1) as reactants. The simulation results are presented in Figure S14. However, regardless of the inclusion of water molecules or the application of different pressure conditions, we do not observe any “C4N2” molecules with the CN-backbone of uracil. We conducted eight more AIMD simulations to investigate the possibility of forming uracil through reactions involving those molecules at various pressures and 1400 K (see the Supporting Information).
@@
To study the ring formation of the uracil molecule, we also selected the “C4N2-1” molecule with large curvature and built a simulation box containing 6 “C4N2-1” and 6 water molecules. Fig. 6 (b) shows that all the “C4N2-1” molecules reacted. At 5.6 GPa and 1400 K, linear “C4N2-1” molecules became 5-membered and 6-membered rings within a shorter time of about 3.6 ps. Particularly, the elements in the six-membered-ring molecule are the same as those in uracil as shown in Fig. 6(a). At 9.3 GPa and 1400 K, the same 6-membered-ring molecule is also formed. These simulations suggest that it is possible for the open-chain form of molecules with same CN-backbone of uracil to transform into the cyclic form. We further performed free energy calculations to analyze this ring formation reaction. We first chose the distance between the head and tail carbon atoms in the open-chain form of uracil as the CV. The enhanced sampling simulation at $\sim$ 10 GPa and 1400 K produced the same 6-membered-ring molecule as in our previous simulation. Note that the free energy profile suggests that the cyclic form is at the metastable state.
@@
Later we chose allophanic acid (C2H4N2O3, number 13 in Fig. 1 ) and (Z)-ethene-1,2-diol (C2H4O2, number 7 in Figure S8) as reactants, and built a new simulation box containing 5 C2H4N2O3, 5 C2H4O2, and 10 H2, as shown in Fig.6 (c). At $\sim$ 10 GPa and 1400 K, a “C5N2” molecule, 3-(3-carboxyureido)-2,3-dihydroxypropanoic acid, was formed at around 28.8 ps, This molecule can be considered as the CN-backbone of thymine (T). As the simulation continued, a “C4N2” molecule, 2,3-dihydroxy-3-ureidopropanoic acid, labeled as “C4N2-2”, was formed. The formation process of these two molecules are shown in Figure S18. We further performed a enhanced sampling simulation to analyze the ring formation process at $\sim$ 10 GPa and 1400 K. The simulation box has one “C4N2-2” molecule and 32 water molecules, as shown in Fig. 6 (c). Finally, the molecule with a uracil ring, known as uracil glycol, was successfully synthesized, which is a uracil-like molecule. The free energy curve in Fig. 6 (c) shows that the cyclic form is more stable than the open-chain form, indicating that this uracil-like molecule is possible to form in geological C-H-O-N fluids at extreme P-T conditions. we do not postulate that uracil molecules cannot be synthesized due to the limited time scale and length scale of such atomistic simulations. As molecules containing the uracil ring have already been observed, we are confident that with a longer duration of chemical reactions, it is highly probable to produce complete uracil molecules in geological fluids. The uracil molecule can then serve as a precursor for the synthesis of other three types of nucleobases, providing the necessary components for forming RNA molecules.
@@ III Conclusions
Extensive AIMD simulations have been performed to investigate the chemical reactions of small inorganic molecules under the extreme conditions found in Earth’s upper mantle. Three corresponding P-T systems were considered, i.e., 10 GPa - 1000 K, 10 GPa - 1400 K, and 13 GPa - 1400 K. The simulation results have revealed the potential formation of substantial quantities of organic matter, with the composition and abundance of CN-containing species being profoundly influenced by the specific P-T conditions. Notably, formamide, recognized as a crucial precursor in the production of biomolecules, is observed to readily form under high P-T conditions, and its formation pathways are compared across different P-T conditions. Furthermore, free energy calculations have confirmed that the 10 GPa - 1400 K condition is particularly favorable for the generation and preservation of CN-containing compounds.
Using the molecules formed, we further explore the formation process of large biomolecules that are directly related to the building blocks of life. Remarkably, glycine, ribose, and uracil-like molecules are successfully synthesized under extreme conditions in our AIMD simulations, thereby providing initial insights into the formation of organic compounds critical to the emergence of life. More importantly, different from that the 6-membered ring of ribose is the majority in solutions at ambient conditions, it is found that formation of 5-membered ring form of ribose is more favorable at Earth’s mantle conditions, which could provide the necessary components for the generation of RNA molecules. Moreover, these compounds have the potential to be transported to the Earth’s surface through geological activities, such as volcanic eruption, consequently contributing to the origins of life. Our comprehensive studies present a novel perspective suggesting that life may have originated from the depths of the Earth’s interior. The findings underscore the profound impact of extreme P-T conditions on the formation of organic matter and provide valuable insights into the mechanisms underlying the emergence of life on our planet.
@@ Methods
III.1 First Principles molecular dynamics simulations (FPMD)
Born-Oppenheimer ab initio molecular dynamics (AIMD) simulations are conducted using the Qbox codegygi2008architecture to investigate the reactions of C-H-O-N fluids under Earth’s upper mantle conditions. The PBE exchange-correlation functional and ONCV pseudopotentials hamann2013optimized are utilized, and the Bussi-Donadio-Parrinello (BDP) thermostat, with a relaxation time of 24.2 fs, is employed to control the temperature bussi2007canonical . To facilitate a larger time step of 0.24 fs, deuterium atoms are used in place of hydrogen atoms in simulations. For the initial setup, the simulation box consists of 15 NH3, 15 H2O, 15 H2, and 15 CO, as depicted in Figure S1. Periodic boundary condition are applied. Three P-T systems are considered based on Earth’s upper mantle conditions: 10 GPa - 1000 K, 10 GPa - 1400 K, and 13 GPa - 1400 K. NPT simulations are carried out initially to achieve the desired system pressure, with the plane-wave energy cutoff set to 85 Ry for the real cell and 70 Ry for the reference cell. The results obtained from NPT simulations are presented in the supporting information (Figure S2). Once the size of the unit cell is converged, a new box will be built based on the equilibrium stage in NPT simulations to run NVT simulations (Table S1). The energy cutoff is adjusted to 65 Ry. Simulations spanning a duration of over 400 ps are conducted under different P-T conditions to ensure the attainment of system equilibrium. All chemical product analyses presented in the main text are obtained from the NVT simulations. Figure S3 shows the total energy and temperature variations during the simulation period. Within the simulation time, the total energy gradually changes little, and the temperature remains stable in all systems. Pressure is also monitored in NVT simulations. To improve computational efficiency, pressure values are not directly imported from NVT simulations but instead calculated by carrying out independent simulations at different time points of NVT simulations with an energy cutoff of 85 Ry. We select random time points under different P-T conditions for the test, which reach approximately 200 ps of NVT simulations, and every data point of pressure is obtained by averaging 12,000 steps, as shown in Figure S4. The total and component values are all within the error ranges. Therefore, we can confidently assert that our NVT simulations are executed under the target P-T conditions.

### v2

@@ II Results and Discussion
We first considered ribose (C5H10O5). The formose reaction is a well-known prebiotic pathway for synthesizing ribose from formaldehyde (CH2O) breslow1959mechanism . Our simulations have generated a few relevant hydrocarbon products (shown in Figures S5 and S6), including formaldehyde molecule (number 1 in Figure S5), which was later chosen as the primary reactant. Additionally, to facilitate the synthesis process, the reactants also included a “C2” molecule, (Z)-ethene-1,2-diol (C2H4O2, number 7 in Figure S5). Initially, we constructed a simulation box consisting of 10 (Z)-ethene-1,2-diol molecules and 5 formaldehyde molecules (Figure S7). After $\sim$76 ps, a “C5” molecule was formed at $\sim$10 GPa and 1400 K. However, it did not transform into the structure of ribose after an extended time. We further tried to add other reactants to the simulation box, and found that after adding 5 water molecules, the ribose molecule appeared, indicating the importance role of water in the formation of ribose. The reaction process is shown in Fig. 5 (a), where the formaldehyde and “C2” molecules react to form a “C3” molecule, which then combines with another “C2” molecule to form a “C5” molecule. This “C5” molecule then evolves into the open-chain form of ribose at $\sim$19 ps.
@@
We calculated the free energy landscape in Fig. 5 (b), where two C-C distances were selected as collective variables labelled as CV1 and CV2. For CV1, two carbon atoms come from formaldehyde and a “C2” molecule, respectively, which are labeled as C1. While for CV2, two carbon atoms are from two different “C2” molecules, marked as C2. The free energy landscape shows a clear energy minimum corresponding to the formation of ribose, but the C-C distance at this minimum is larger than a typical C-C bond distance. We examined the biased molecular dynamics trajectories and found that the carbon backbone of the generated ribose molecule has the C1-C2-C1-C2 structure. By measuring the distances of C1-C1 and C2-C2 , we confirmed that they are consistent to the distances observed in the free energy landscape. We also randomly selected another initial configuration to confirm that the ribose molecule is thermodynamically stable, as shown in Figure S8.
@@
Previous studies showed that the cyclic forms of ribose are abundant in aqueous solutions at room temperature drew199813c . However, the cyclic forms of ribose were not detected in our simulations (Fig. 5 (a)). The limited size of our simulation boxes may restrict the configuration changes of ribose, thereby decreasing the likelihood of the C4’ or C5’ hydroxyl group attacking the C1’ aldehyde group. This limitation is unfavorable for the production of the furanose form (five-membered ring) or pyranose form (six-membered ring). To further study the cyclic formation mechanism, we carried a few additional simulations and finally chose the curved open-chain ribose molecules obtained from our previous simulations, as it appears that the C4’ or C5’ hydroxyl group has a greater chance of reacting with the C1’ aldehyde group under pressure. We built a new simulation box containing three of these ribose molecules together with 18 water molecules. After about 36 ps, we found the direct formation of a furanose form of ribose. Fig. 5 (c) shows the mechanism of ring formation. In this process, the C4’ hydroxyl group first loses a hydrogen atom, leaving a carbonyl group. The oxygen atom then attacks C1’ to form the ring, and finally, the carbonyl group on C1’ receives a dissolved proton in the solution to generate the five-membered ribose ring. Interestingly, no six-membered ring was found in our simulations at 10 GPa and 1400 K.
@@
We further applied free energy calculations to study the formation ability of five- and six-membered ribose rings at $\sim$ 10 GPa and 1400 K. We chose two C-O distances as the CVs for the two rings, as shown in Fig. 5 (d). The energy minima are at the C-O distance of 1.37 Å for the two ribose forms, indicating the formation of C-O bonds and the rings. Besides, the formation of the C-O bonds lowers the free energy by 11.57 kcal/mol for the five-membered ribose ring and 5.85 kcal/mol for the six-membered ribose ring, indicating that the ring formation is thermodynamically more stable and the furanose form of ribose is more favored than the pyranose form at $\sim$ 10 GPa and 1400 K. The free energy result explained why the the pyranose form of ribose is absent in our AIMD simulations. The spacial limitation under high pressure may hinder the formation of larger cyclic forms of molecules, so the energy barrier of forming a five-membered ring is lower than that of a six-membered ring.
@@
Note that at ambient conditions, the naturally-occurring ribose exists as a mixture of cyclic forms in equilibrium with its open-chain form in aqueous solutions, and the pyranose form of ribose is the predominant conformation, accounting for approximately 70% dewick2013essentials ; dass2021equilibrium . However, despite the thermodynamic stability of the pyranose form of ribose over the furanose form at ambient conditions, the exclusive constituent of the carbohydrate backbone of RNA remains the furanose form of ribose. The reason behind it is still unknown. There are several hypotheses proposed to explain this preference, including silicate(borate)/ribose complexes kolb2004complexes ; lambert2004silicate and temperature gradients dass2021equilibrium . Our finding suggests that extreme P-T conditions may also generate the necessary component for assembling RNA molecules. Consequently, it highlights the potential role of upper mantle environment in facilitating the emergence and development of RNA-based life.
@@
We then shifted our attention to the chemical evolution of nucleobases, which are another essential building blocks of genetic information. There are five naturally-occurring nucleobases: adenine (A), cytosine (C), guanine (G), thymine (T), and uracil (U). Of these, uracil is unique to RNA, so our studies primarily focused on investigating the formation of uracil molecules under extreme conditions. We first studied the formation of the open-chain “C4N2” molecule which has the same CN-backbone as the open-chain form of uracil, as shown in Fig. 6 (a). Previous experimental studies on the origins of life have suggested that formamide is a crucial molecule for the prebiotic formation of nucleobases, in the presence of mineral or metal oxide catalysts saladino2012formamide , heatingenchev2021chemical , and high-energy free radicals ferus2015high ; ferus2014high , so here we also first chose formamide as a reactant. However, after about 144 ps, we did not observe the formation of the “C4N2” molecule at 10 GPa and 1400 K, suggesting that it may be difficult for small molecules to directly synthesize uracil in the AIMD simulations (Figure S9). Later, we chose the chemical reactants that have closer structural similarities to the “C4N2” molecule. The reactant molecules, including “C4N”, “C3N2”, “C2N2”, and “C3”, were generated from our first-stage AIMD simulations. A recent study suggested that uracil could be synthesized from the reactions between urea and “C3” molecules choe2022mechanism , so we specifically selected the C3H4O4 molecule (number 10 in Figure S5) and the urea molecule (CH4N2O, number 11 in Fig. 1) as reactants. The simulation results are presented in Figure S10. However, regardless of the inclusion of water molecules or the application of different pressure conditions, we did not observe any “C4N2” molecules with the CN-backbone of the open-chain form of uracil. Then we tried to find another pathways and conducted six more AIMD simulations to investigate the possibility of forming uracil through reactions involving above-mentioned molecules at various pressures and 1400 K (see Figures S11, S12 and S14).
@@
To study the ring formation of uracil, we also selected the “C4N2-1” molecule with large curvature and built a simulation box containing 6 “C4N2-1” and 6 water molecules. Fig. 6 (b) shows that all the “C4N2-1” molecules reacted. At 5.6 GPa and 1400 K, open-chain “C4N2-1” molecules became five-membered and six-membered rings within a shorter time of about 3.6 ps. Particularly, the elements in the six-membered-ring molecule are the same as those in uracil as shown in Fig. 6(a). At 9.3 GPa and 1400 K, the same six-membered-ring molecule was also formed. These simulations suggest that it is possible for the open-chain form of uracil to transform into the cyclic form. We further performed free energy calculations to analyze this ring formation reaction. We first chose the distance between the head and tail carbon atoms in the open-chain form of uracil as the CV. The enhanced sampling simulation at $\sim$ 10 GPa and 1400 K produced the same six-membered-ring molecule as in our unbiased simulations. Note that the free energy profile suggests that the cyclic form is in a metastable state, as shown in Fig. 6(b).
@@
Later we chose allophanic acid (C2H4N2O3, number 13 in Fig. 1 ) and (Z)-ethene-1,2-diol (C2H4O2, number 7 in Figure S5) as reactants, and built a new simulation box containing 5 C2H4N2O3, 5 C2H4O2, and 10 H2, as shown in Fig.6 (c). At $\sim$ 10 GPa and 1400 K, a “C5N2” molecule, 3-(3-carboxyureido)-2,3-dihydroxypropanoic acid, was formed at around 28.8 ps, This molecule has the same CN-backbone as the open-chain form of thymine (T). As the simulation continued, a “C4N2” molecule, 2,3-dihydroxy-3-ureidopropanoic acid, labeled as “C4N2-2”, was formed, which has the same CN-backbone as the open-chain form of uracil. The formation process of these two kinds of molecules are shown in Figure S15. We further performed a enhanced sampling simulation to analyze the ring formation process at $\sim$ 10 GPa and 1400 K. The simulation box has one “C4N2-2” molecule and 32 water molecules, as shown in Fig. 6 (c). Finally, a uracil-like molecule, called uracil glycol, was synthesized, which has a uracil ring. The free energy curve in Fig. 6 (c) shows that the cyclic form is more stable than the open-chain form, indicating that the uracil-like molecule is possible to form in geological C-H-O-N fluids at extreme P-T conditions. The uracil-like molecule can then serve as a precursor for the synthesis of other three types of nucleobases, providing the necessary components for forming RNA molecules.
Geochemical evidence of biotic activity suggests that the emergence of life can be traced back to as early as 3.8 billion years ago mojzsis1996evidence . The modern style of plate tectonics may have started about 3.2 billion years ago brenner2020paleomagnetic , but some form of plate tectonics may have already been active around 4.0 billion years ago. and these tectonic activities might carry water, carbon, and other volatile substances down into the mantle korenaga2013initiation ; HOLDER2024 ; blichert2008hafnium ; hopkins2008low ; huang2022barium . Because of reducing environments in early Earth’s mantle aulbach2016evidence , oxidized carbon and nitrogen can be reduced to CO and NH3/NH+${}_{4}^{+}$, respectively holm2009reduction (see Figure S16), and they may further synthesize to become key building blocks of life, which may reach the surface of early Earth. Our findings also have important implications for organic molecules in interstellar space. The reactant molecules H2O, H2, CO and NH3 largely exist in interstellar clouds and circumstellar shells, together with more than 200 complex organic species, such as aldehydes, alcohols, acids, amines and carboxamides guelin2022organic . It is unclear how those complex organic species are generated. Our current study suggests that the extreme pressure and temperature inside giant planets and their moons neri2020carbonaceous ; nettelmann2016uranus , like Jupiter and Saturn, may help to synthesize those complex organic molecules postberg2018macromolecular ; waite2009liquid ; tosi2024salts .
@@ III Conclusions
Extensive AIMD simulations have been performed to investigate the chemical reactions of small proto-organic molecules under the extreme P-T conditions. Three corresponding P-T systems were considered, i.e., 10 GPa - 1000 K, 10 GPa - 1400 K, and 13 GPa - 1400 K. The simulation results have revealed the potential formation of substantial quantities of organic matter without any catalysts, with the composition and abundance of CN-containing species being profoundly influenced by the specific P-T conditions. Notably, formamide, recognized as a crucial precursor in the production of biomolecules, is observed to readily form under extreme P-T conditions, and its formation pathways are compared across different P-T conditions. Furthermore, free energy calculations have confirmed that the 10 GPa - 1400 K condition is particularly favorable for the generation and preservation of CN-containing compounds.
Using the molecules formed, we have further explored the formation process of large biomolecules that are directly related to the building blocks of life. Remarkably, glycine, ribose, urea, and uracil-like molecules are successfully synthesized under extreme conditions in our AIMD simulations, thereby providing initial insights into the formation of organic compounds critical to the emergence of life. More importantly, different from that the six-membered ring of ribose is the majority in solutions at ambient conditions, it is found that formation of five-membered ring form of ribose is more favored at extreme P-T conditions, which could provide the necessary components for the generation of RNA molecules. Moreover, these compounds have the potential to be transported to Earth’s surface through geological activities, such as mantle convection, consequently contributing to the origins of life. The complex organic species found in interstellar space may be also related to the chemical processes inside giant planets and their moons at extreme P-T conditions. Our comprehensive studies present a novel perspective extending the cradle of life planetary interiors. The findings underscore the profound impact of extreme P-T conditions on the formation of organic matter and provide valuable insights into the mechanisms underlying the emergence of life within planetary bodies.
@@ Methods
III.1 Ab initio molecular dynamics simulations (AIMD)
We carried out AIMD simulations with the Born-Oppenheimer approximation using the Qbox code gygi2008architecture . We applied the PBE exchange-correlation functional in density functional theory perdew1996generalized . It is known that the PBE functional may be insufficient to describe aqueous solutions at ambient conditions, but our previous studies found that it works better at extreme P-T conditions than at ambient conditions pan2013dielectric ; pan2016fate ; stolte2019large ; stolte2021water , and many other studies also showed that this semi-local functional is reliable to study matter under extreme conditions spanu2011stability ; cheng2023thermodynamics ; zhang2023pbe . We used the ONCV pseudopotentials hamann2013optimized , and the Bussi-Donadio-Parrinello (BDP) thermostat, with a relaxation time of 24.2 fs, to control the temperature bussi2007canonical . To facilitate a larger time step of 0.24 fs, deuterium atoms were used in place of hydrogen atoms in our simulations. The simulation boxes have periodic boundary conditions. We first performed NPT simulations to achieve the desired pressures, configuring the plane-wave energy cutoff to 85 Ry and the energy cutoff for stress confinement potential to 70 Ry. We applied the simulation box sizes (Table S1) obtained from NPT simulations in the following NVT simulations for production runs, where the plane-wave energy cutoff was 65 Ry. During NVT simulations, we selected random time points to carry out independent simulations with energy cutoff of 85 Ry to monitor pressure (Figure S2). Our total AIMD simulation time exceeds 2.5 ns.